OpenStudy (winsteria):

SOMEONE PLEASE HELP ME IF YOU DO I'LL GIVE YOU A MEDAL AND FAN What is the reciprocal of the number? Column A: 1. -3/7 2. 3/7 3. -7/3 4. 7/3 Column B: A. -7/3 B. -3/7 C. 3/7 D. 7/3 E. 1 (YOU HAVE TO MATCH UP THE OPTIONS FROM EACH COLUMN TO EACHOTHER)

OpenStudy (winsteria):

@sangya21 @TheSmartOne

OpenStudy (yousir):

you called?

OpenStudy (winsteria):

what?

OpenStudy (yousir):

thesmartone... get it

OpenStudy (winsteria):

ohh

OpenStudy (anonymous):

Reciprocal of a = 1/a so what will be the answers?

OpenStudy (yousir):

ow thx for stealin my thundre

OpenStudy (winsteria):

Uhhh

OpenStudy (winsteria):

1 is C?

OpenStudy (anonymous):

No cause negative sign is missing

OpenStudy (winsteria):

So 1 is B?

OpenStudy (anonymous):

reciprocal of -3/7 = 1/(-3/7) = -7/3

OpenStudy (winsteria):

ohhhh

OpenStudy (winsteria):

So 2 is D?

OpenStudy (anonymous):

correct

OpenStudy (winsteria):

And 3 is B? And 4 is C?

OpenStudy (anonymous):

correct. :)

OpenStudy (winsteria):

oh okay, yay thank you so much! :)

OpenStudy (anonymous):

Anytime :)